# Literature Review: Cyber Security Monitoring in Maritime

### v1

@@ Abstract
In recent years, many cyber incidents have happened in the maritime sector, targeting the information technology (IT) and operational technology (OT) infrastructure. Although several systematization-of-knowledge papers have been published in the maritime field, none of the previous studies has focused on cyber security monitoring, which aims at timely detection of cyber attacks with automated methods.
The current article addresses this research gap and surveys the methods, algorithms, tools and architectures used for cyber security monitoring in the maritime sector. For the survey, a systematic literature review of cyber security monitoring studies is conducted in this article, following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) protocol.
The first contribution of this article is the bibliometric analysis of related literature and the identification of the main research themes in previous works. For that purpose, our article presents a taxonomy for existing studies which highlights the main properties of maritime cyber security monitoring research. The second contribution of this article is an in-depth analysis of previous works and the identification of research gaps and limitations in existing literature. Based on our findings, we outline future research directions for cyber security monitoring in the maritime field.
@@
cyber security, cyber security monitoring, intrusion detection, maritime, monitoring, literature review
@@ I Introduction
Cyber attack detection and mitigation is further complicated by unique features of maritime systems (see [13, 14, 15] for a more detailed discussion). Maritime systems usually consist from a number of different IT and OT networks that connect a wide variety of maritime devices which communicate over specialised protocols such as NMEA. Therefore, the detection of attacks conducted in such networks requires detailed analysis of these maritime-specific protocols by dedicated solutions. Furthermore, ships have a limited network connectivity with the shore, which leads to the need for autonomous attack detection functionality on vessels. Also, space limitations and other physical constraints do not often allow to operate significant computational resources (e.g., server clusters) on board of ships, introducing the need for lightweight and resource-efficient cyber security solutions.
@@
Although security monitoring technologies have been widely adopted by traditional (i.e., non-maritime) organisations and much academic research has been conducted in this field, security monitoring remains understudied in the maritime domain. For example, detailed cyber security guidelines by BIMCO [12] touch security monitoring only briefly, mentioning the use of network IDS and malware detection for security monitoring purposes, but failing to provide more detailed implementation recommendations to build maritime SOCs and other maritime security monitoring solutions. As another example, cyber risk management guidelines by IMO [18] do not address cyber security monitoring at all. As pointed out in [13], the generic nature of existing guidelines complicates the creation of maritime security monitoring systems, and dedicated research in this domain is needed. Also, as discussed in Section IV of this paper, maritime security monitoring is a relatively new research field, with most academic papers being published in the last 5-6 years. Due to the novelty of this research field, no domain overview studies have been published in maritime security monitoring that would analyse existing academic literature and identify the main research topics and open challenges.
This article fills the aforementioned research gap and specifically targets the research on security monitoring in the maritime domain. The purpose of the current study is to provide a Systematic Literature Review (SLR) of research papers published in the field. Note that our study focuses on peer-reviewed academic papers, excluding grey literature from consideration. A similar approach has been used in recent SLRs that address maritime cyber security, risk, and safety [3, 19]. Using the same approach allows us to evaluate existing literature from an academic perspective and to identify research themes and gaps in analysed papers.
Apart from considering peer-reviewed publications only, we did not impose any additional limitations to the literature selection process. As a result, analysed papers cover a wide variety of maritime cyber security monitoring topics, including the design of a maritime SOC, the creation of testbed environments for maritime security monitoring experiments, and various security monitoring methods (e.g., ML-based approaches) for navigational systems, vessel CPSs, maritime IoT networks, etc.
@@
The research gaps we have found are the following. First, existing research has not utilised publicly available maritime-specific datasets. Also, many existing papers are suffering from several evaluation issues, for example, failing to employ standard performance metrics, to measure the computational cost of the proposed methods, and to release the experiment code publicly. These deficiencies are in conflict with the best practices of the security monitoring research, decreasing the reproducibility of the experiments and complicating the assessment of the research results [17, 20]. Furthermore, we discovered a number of unexplored research topics during literature analysis. For example, existing studies on maritime cyber security monitoring have not addressed advanced ML issues like concept drift and learning models for resource-constrained environments. As another example, human aspects of maritime security monitoring have received little attention in the existing literature. For a detailed discussion of identified research gaps, please see Section VI.
The remainder of this article is organised as follows. Section II describes similar domain overview and SLR papers, Section III formulates the research questions of our study and discusses the selection criteria for research covered by our study, and Section IV contains the bibliometric analysis of the research papers selected for the study. Section V presents the taxonomy of relevant research and provides an in-depth analysis of selected research papers. Section VI discusses open issues and research gaps in the existing research literature, and Section VII concludes the article.
@@ II Related work
The purpose of this section is to provide an overview of similar SLR papers in the maritime domain. At the end of this section, we will also mention some systematisation-of-knowledge papers on intrusion and anomaly detection in non-maritime environments. The maritime SLR papers discussed in this section are summarised in Table I. As can be seen from Table I, existing SLRs have mainly focused on the analysis of past cyber incidents and ways for mitigating cyber threats in the future. For that purpose, the authors have proposed various measures, most notably by changing existing regulations and policies, and by improving the cyber security training and cyber awareness of personnel. Only one SLR paper [3] covers some security monitoring studies, but their treatment has remained very brief without a deeper analysis. Although the papers from Table I do not focus on maritime security monitoring, we will discuss these papers for the sake of the completeness of our study, and to illustrate that the treatment of maritime security monitoring has remained too brief in existing SLRs.
TABLE I: Maritime Cyber Security SLR Papers
@@
[22] 2023 cyber risk management, analysis of regulations and policies missing
@@
[2] 2022 ensuring cyber security during digitalisation of a marine company missing
[5] 2022 analysis of cyber threats, regulations and policies missing
@@
[10] 2020 analysis of attacks against navigational systems and mitigation measures missing
@@
Yu et al. [21] conducted an SLR on cyber security in the maritime domain, describing common trends in related literature. As commonly occurring research themes, the authors identified the impact assessment for maritime cyber threats, recognition of maritime cyber threats (e.g., through vulnerability management tools or through interviews with human personnel), and categorisation and mitigation of maritime cyber risks. The study also presented recommendations for risk mitigation, and discussed gaps in existing literature and avenues for further research. As a supplementary material to their paper, the authors published a list of 21 maritime cyber attacks that have received media coverage and public attention, providing a short description of each cyber attack. Although the study briefly mentioned couple of security monitoring approaches for cyber risk mitigation (e.g., the use of ML methods for AIS anomaly detection), these approaches were not elaborated further.
Androjna et al. [10] investigated the prospect of cyber threats and confirmed incidents within the maritime industry, aiming to identify the emerging cyber trends and challenges impacting safe navigation and marine shipping. The authors focused especially on the vulnerabilities of the Global Positioning System (GPS) as one of the core components for safe navigation. An SLR was conducted supplemented by an analysis of the aforementioned specific spoofing event that occurred on Elba Island in 2019. The authors summarised that a change in thinking is important to drive more processes towards cyber security as well as the need for manufacturers to enhance the cyber security of their designs. Encouraged by ongoing national defence research on alternative positioning, navigation and timing systems, as well as by IMO’s e-Navigation concept, the authors also proposed the use of systems that offer different positioning, navigation, and timing capabilities aboard ships, serving as to facilitate GPS-inclusive navigation. Furthermore, the authors recommended utilising various satellite navigation constellations, both public and private, alongside earthbound sub-systems of an enhanced LOng-RAnge Navigation (eLORAN) system as well as advanced ship aiding systems at ports. This approach aims to facilitate the maritime domain with the required improved defense against cyber threats in the near future.
Kechagias et al. [2] reviewed the aspects that should be considered during the digitalisation of a marine company. The review was focused on cyber security within the maritime domain, presenting an observation of the current state of the topic. The authors described how real-world organisations can manage and develop overall consciousness and intelligence in safeguarding shipping from malicious actors. The significance and main attributes of cyber security in the maritime domain were introduced and analysed. The authors outlined systemic methodologies by pointing out improved procedures and policies. Initially the plan-do-check-act (PDCA) method was introduced for emphasising continual enhancement. Another method then emphasised the effective management of the three primary aspects of security i.e. procedures, the human factor, and technology. Additionally, to augment comprehension of the topic, an audit survey of vessel cyber security was conducted.
Schinas and Metzgera [22] discussed maritime cyber risks, the IMO instruments, the trend of digitalisation, and insurance-related literature. The authors revisited the concept of seaworthiness and examined it from the perspective of cyber risks with a special focus on maritime autonomous surface ships, figuring out the importance of the regulations’ update. As a response to the identified policy gap, the authors introduced the term “cyber-seaworthiness”.
@@
In their paper, Bronk and deWitte [5] described cyber issues of maritime vessels and ports, providing examples of critical shipboard navigational systems and vulnerable software solutions used in ports. The authors also reviewed some maritime related cyber security policies and guidelines. For example, the US Coast Guard Cyber Strategy emphasises the importance of risk management which relies on promoting cyber risk awareness and reduction of vulnerabilities. Also, IMO’s cyber risk management guidance describes eight areas which may contain vulnerable systems such as bridge systems, cargo handling systems, public networks designed for serving passengers, etc. The authors concluded by identifying research on protecting the IT systems of ships and ports as an important future research direction.
@@
Another SLR by Chaal et al. [19] analysed the research literature on the risk, reliability, and safety of autonomous ships. Whereas cyber security was not the primary focus of this SLR, it was identified as an important part of the ship’s safety. Also, the study indicated that the cyber security of autonomous vessels has started to receive more attention from researchers recently. However, the study did not include the treatment of cyber security monitoring topics.
Another study by Ben Farah et al. reviewed cyber incidents in the maritime domain and provided a relatively concise bibliographic analysis [4]. The authors provided an extensive elaboration of maritime sub-systems located in vessels and ports, in addition to a review of the potential cyber attacks targeting these systems. Similarly to previous studies described in this section, the authors did not focus on cyber security monitoring issues.
Intrusion and anomaly detection have been broadly addressed in various review studies without focusing on specific systems or network types [23, 24, 25]. Following the introduction of the Internet of Things (IoT) networks, a huge body of literature has researched methods and approaches to run intrusion detection functions within the limits of this network category. Some papers reviewed the relevant literature and discussed the challenges and solutions for IoT environments [26, 27]. Vehicular ad hoc networks (VANETs) and medical systems are also other important system categories that have recently been the subject of various literature review papers in the intrusion detection domain [28, 29].
Our article addresses a gap in the existing SLR literature on the cyber security of the maritime sector. Unlike previous analyses that studied various aspects, our contribution specifically delves into maritime security monitoring systems, an area that has not been adequately covered thus far.
@@ III Methodology
In this section, the systematic review process will be explicated, describing the undertaken methodological approach together with the inclusion and exclusion criteria for relevant research, as well as the overarching review objectives.
@@ III-A Aims and research questions
RQ1: What are the statistical characteristics of the papers that constitute the relevant literature? Analysing the literature through bibliometric methods will provide insights on the publication timeline, the types of venues in which the works were presented, the most prolific authors, and other properties of related work. RQ1 is answered by the discussion in Section IV.
@@ III-B1 Eligibility Criteria
Initially, criteria need to be set that guide the search process and help to filter out irrelevant works. In this article, papers written in English and published in peer-reviewed journals and conference/workshop proceedings are included. When searching for relevant papers, we did not set an age limit for publications (i.e., we did not exclude papers published before a specific date). Papers that did not focus on maritime cyber security monitoring were excluded, just like papers focusing on non-IT-related methods from the fields of electrical engineering, physics, and other fields not related to IT. Editorials, opinions, keynotes, abstracts, tutorial summaries, position papers, panel discussions, technical reports and posters were also excluded. The duplicated publications because of different versions were also excluded, with only the most recent version being included. Since this SLR considers peer-reviewed academic papers, grey literature was excluded from this research.
Inclusion and exclusion criteria are summarised below. For a paper to be included, all inclusion criteria described in Section III-B1a must be met. On the other hand, a paper is excluded if it matches any of the exclusion criteria from Section III-B1b.
III-B1a Inclusion Criteria
@@
  • Studies in the domain of cyber security in the maritime sector with a focus on monitoring and intrusion detection.
III-B1b Exclusion Criteria
  • Studies that do not explicitly consider cyber security monitoring and intrusion detection.
@@ III-B2 Search Strategy
III-B2a Search for Related Publications
@@
III-B2b Keyword Selection
The selection of keywords heavily influences the results of both automatic and manual searches. During the search phase, several keywords and their combinations were tried on search engines for electronic libraries.
@@
All the search engines that we used support the creation of search queries with Boolean operators (AND, OR and NOT), which helps to narrow the focus of the search and limit the number of hits. Also, the engines allow to apply the search expression on the title, abstract, keywords, metadata, etc. of the papers. The time frame of the search is also definable, and the wildcard characters (such as asterisk (*) and question mark (?)) allow further flexibility during the search.
@@
In the first phase of keyword selection, the keywords related to the maritime, like marine, ship, and navy were selected to define the sector. These keywords combined with the “OR” operator led to papers meant to protect areas against marine infrastructure. Then, the focus was specified —in the same way— by keywords and phrases, like monitor, “intrusion detection”, IDS, “attack detection”, etc. The query made of these keywords brought many publications, but the majority were irrelevant because of the territorial defence mechanisms and different anomalies in the sector. By adding “cyber” to these keywords, the focus was narrowed down sufficiently enough.
III-B2c Automated Search on Bibliographical Sources
@@
During the initial stage of the automated search on bibliographical sources several trials were applied, using different combinations of keywords according to the main objectives of the review, with the aim to construct an optimal keyword set. The final keyword set was then used to construct the search query to be instantiated and executed in various search engines of electronic libraries. The selected electronic libraries include the most important conference proceedings and highest impact journals. The following electronic libraries were used with the same query applied after it was adapted accordingly:
  • ACM digital library (http://portal.acm.org/)
    [[Abstract: monitor*] OR [Abstract: “intrusion detection”] OR [Abstract: “anomaly detection”] OR [Abstract: “attack detection”] OR [Abstract: ids]] AND [Abstract: maritime ship navy navy sea marine ais]
  • IEEE Xplore (http://ieeexplore.ieee.org/Xplore/)
    (“All Metadata”: maritime OR “All Metadata”: ship OR “All Metadata”: navy OR “All Metadata”: sea OR “All Metadata”: marine) AND (“All Metadata”: monitor* OR “All Metadata”: “intrusion detection” OR “All Metadata”: “anomaly detection” OR “All Metadata”: “attack detection” OR “All Metadata”: IDS) AND (“All Metadata”: cyber)
  • Elsevier Scopus (https://www.scopus.com/)
    (TITLE-ABS-KEY ( maritime OR navy OR ship OR marine OR sea OR ais) AND TITLE-ABS-KEY (“intrusion detection” OR monitoring OR “anomaly detection” OR “attack detection” OR ids) AND TITLE-ABS-KEY(cyber))
  • Web of Science (https://www.webofscience.com/)
    cyber (All Fields) AND “intrusion detection” OR monitoring OR “anomaly detection” OR “attack detection” OR ids (All Fields) AND maritime OR navy OR ship OR marine OR sea OR ais (All Fields)
III-B2d Paper Selection Process
@@
The automated search on bibliographical sources resulted in over 1000 papers. After the library search, the four co-authors processed one library each in the form of independent work, and the findings were added to a shared database that helped to indicate the duplicated items. The next step in the selection of relevant papers was the analysis of the title and the abstract. If a paper was irrelevant based on the title and/or the abstract, it was excluded. (In case of doubt, the researcher also read the full text.) If title and abstract were in the scope, the paper became a candidate. Since one more co-author joined to keep the workload balanced, the candidate papers were re-shared among the authors, who reread the full text.
@@
If a paper remained relevant, its main characteristics were noted to support our findings. After the full-text reading of the candidates, we had a set of papers conforming to the inclusion criteria. That constituted the base of the backward snowballing procedure, i.e., additional search by manually checking and analysing the related references. Forward snowballing was also applied to find relevant papers lately published citing any of the selected papers. This additional task ensured the inclusion of every relevant study [33]. The eligibility analysis ended up in 23 relevant papers. At this point, the bibliometric analysis was written, and the methodology for analysis was developed.

### v2

@@ Abstract
In recent years, many cyber incidents have occurred in the maritime sector, targeting the information technology (IT) and operational technology (OT) infrastructure. Although several literature review papers have been published in the maritime field, none of the previous studies has focused on cyber security monitoring, which aims at timely detection of cyber attacks with automated methods. The current article addresses this research gap and surveys the methods, algorithms, tools and architectures used for cyber security monitoring in the maritime sector. For the survey, a systematic literature review of cyber security monitoring studies is conducted. The first contribution of this article is the bibliometric analysis of related literature and the identification of the main research themes in previous works. For that purpose, our article presents a taxonomy for existing studies which highlights the main properties of maritime cyber security monitoring research. The second contribution of this article is an in-depth analysis of previous works and the identification of research gaps and limitations in existing literature. Based on our findings, we outline future research directions for cyber security monitoring in the maritime field.
@@
cyber security, cyber security monitoring, intrusion detection, literature review, maritime, monitoring
@@ I Introduction
Cyber attack detection and mitigation is further complicated by unique features of maritime systems (see [13, 14, 15] for a more detailed discussion). Maritime systems usually consist from a number of different IT and OT networks that connect a wide variety of maritime devices which communicate over specialised protocols such as NMEA. Therefore, the detection of attacks conducted in such networks requires detailed analysis of these maritime-specific protocols by dedicated solutions. Furthermore, ships have a limited network connectivity with the shore, which leads to the need for autonomous attack detection functionality on board. Also, space limitations and other physical constraints do not often allow to operate significant computational resources (e.g., server clusters) on board of ships, introducing the need for lightweight and resource-efficient cyber security solutions.
@@
Although security monitoring technologies have been widely adopted by traditional (i.e., non-maritime) organisations and much academic research has been conducted in this field, security monitoring remains understudied in the maritime domain. For example, detailed cyber security guidelines by BIMCO [12] touch security monitoring only briefly, mentioning the use of network IDS and malware detection for security monitoring purposes, but failing to provide more detailed implementation recommendations to build maritime SOCs and other maritime security monitoring solutions. As another example, cyber risk management guidelines by IMO [18] do not address cyber security monitoring at all. As pointed out in [13], the generic nature of existing guidelines complicates the creation of maritime security monitoring systems, and dedicated research in this domain is needed. Also, as discussed in Section IV of this paper, maritime security monitoring is a relatively new research field. Due to the novelty of this research area, no domain overview studies have been published in maritime security monitoring that would analyse existing academic literature and identify the main research topics and open challenges.
This article fills the aforementioned research gap and specifically targets the research on security monitoring in the maritime domain. The purpose of the current study is to provide a Systematic Literature Review (SLR) of research papers published in the field. Note that our study focuses on peer-reviewed academic papers, excluding grey literature from consideration. A similar approach has been used in several recent SLRs that address maritime cyber security, risk, and safety [3, 19, 20]. Using the same approach allows us to evaluate existing literature from an academic perspective and to identify research themes and gaps in analysed papers.
@@
The remainder of this article is organised as follows. Section II describes similar literature review papers, Section III formulates the research questions of our study and discusses the selection criteria for research covered by our study, and Section IV contains the bibliometric analysis of the research papers selected for the study. Section V presents the taxonomy of relevant research and provides an in-depth analysis of selected research papers. Section VI discusses open issues and research gaps in the existing research literature, and Section VII concludes the article.
@@ II Related work
The purpose of this section is to provide an overview of similar literature review papers in the maritime domain (see Table I). According to Table I, existing papers have mainly focused on the analysis of past cyber incidents and ways for mitigating cyber threats in the future. For that purpose, the authors have proposed various measures, most notably by changing existing regulations and policies, and by improving the cyber security training and cyber awareness of personnel. Only one paper [3] covers some security monitoring studies, but their treatment has remained very brief without a deeper analysis. Although the papers from Table I do not focus on maritime security monitoring, we will discuss these papers for the sake of the completeness of our study, and to illustrate that the treatment of maritime security monitoring has remained too brief in existing literature reviews.
TABLE I: Maritime cyber security literature review papers
@@
[19] 2024 cyber attacks against MSCN and practices for prevention and mitigation missing
[21] 2024 cyber attacks on autonomous vessels and attack mitigation techniques missing
[20] 2024 cyber threat modelling and risk assessment in the field of maritime missing
@@
Clavijo Mesa et al. [19] conducted an SLR on cyber attacks in maritime supply chain networks (MSCN), providing recommendations on the prevention and mitigation practices. The paper introduced a comprehensive taxonomy for cyber attacks which categorised them based on a number of factors, including the hacker group, attack type, affected systems, geopolitical impact, etc. According to the study, the key cyber security threats to MCSN include malware, Distributed Denial of Service (DDoS) and Denial of Service (DoS) attacks. For prevention and mitigation, the paper proposed a number of technical and policy-level measures (e.g., the use of multi-factor authentication, standardisation of cyber security practices, etc.), and practices for increasing cyber security awareness (e.g., development of training courses about attack prevention). The study briefly mentioned some security monitoring technologies without a deeper analysis.
Symes et al. [21] reviewed the literature focusing on the cyber attacks on autonomous vessels. The paper provided an overview on cyber attack types and described a number of well-known past security incidents in the maritime sector. The paper also discussed possible cyber attack mitigation techniques, mentioning some commonly used security monitoring techniques like network traffic monitoring without a more detailed treatment. For protecting autonomous vessels from cyber attacks, the study provided a number of recommendations, including employee training, regular updating and patching of systems, the use of strong passwords and multi-factor authentication, etc.
@@
Yu et al. [22] reviewed the papers on cyber security in the maritime domain, describing common trends in relevant literature. As commonly occurring research themes, the authors identified the impact assessment for maritime cyber threats, recognition of maritime cyber threats (e.g., through vulnerability management tools or through interviews with human personnel), and categorisation and mitigation of maritime cyber risks. The study also presented recommendations for risk mitigation, and discussed gaps in existing literature and avenues for further research. As a supplementary material to their paper, the authors published a list of 21 maritime cyber attacks that have received media coverage and public attention, providing a short description of each cyber attack. Although the study briefly mentioned couple of security monitoring approaches for cyber risk mitigation (e.g., the use of ML methods for AIS anomaly detection), these approaches were not elaborated further.
@@
Erbas et al. [20] performed an SLR that critically examined and compared existing threat modelling and risk assessment methods in ship cyber security, developing a taxonomy for them. The authors analysed 25 scientific papers to understand the evolving landscape of cyber security practices for manned and autonomous ships. Significant inconsistencies were observed in current approaches and key challenges were identified, underscoring the urgent need for standardised threat modelling and risk assessment frameworks that consider also the special requirements of autonomous ships. Furthermore, the authors highlighted disparities, as well as inadequacy in tool support for the reviewed threat modelling and risk assessment methods. The need for more systematic validation of expert knowledge was also deemed critical. The topic reviewed by Erbas et al. is complementary to security monitoring, as threat modelling and risk assessment can provide direct inputs for enhanced monitoring solutions.
@@
A bibliometric review by Chaal et al. [23] analysed the research literature on the risk, reliability, and safety of autonomous ships. Whereas cyber security was not the primary focus of this paper, it was identified as an important part of the ship’s safety. Also, the study indicated that the cyber security of autonomous vessels has started to receive more attention from researchers recently. However, the paper did not include the treatment of cyber security monitoring topics.
Another study by Ben Farah et al. reviewed cyber incidents in the maritime domain and provided a concise bibliographic analysis [4]. The authors provided an extensive elaboration of maritime sub-systems located in vessels and ports, in addition to a review of the potential cyber attacks targeting these systems. Similarly to previous studies described in this section, the authors did not focus on cyber security monitoring issues.
Our article addresses a gap in the existing SLR literature on the cyber security of the maritime sector. Unlike previous analyses that have largely focused on the cyber threats and risks, our contribution specifically delves into maritime security monitoring systems, an area that has not been adequately covered thus far.
@@ III Methodology
In this section, the systematic literature review process will be discussed, describing the undertaken methodological approach together with the inclusion and exclusion criteria for relevant research, as well as the overarching review objectives.
@@ III-A Aims and research questions
RQ1: What are the statistical characteristics of the papers that constitute the relevant literature? Analysing the literature through bibliometric methods will provide insights on the publication timeline, the types of venues in which the works were presented, and other properties of related work. RQ1 is answered by the discussion in Section IV.
@@ III-B1 Eligibility Criteria
Initially, criteria need to be set that guide the search process and help to filter out irrelevant works. In this article, papers written in English and published in peer-reviewed journals and conference/workshop proceedings were included. When searching for relevant papers, we did not set an age limit for publications (i.e., we did not exclude papers published before a specific date). Papers that did not focus on maritime cyber security monitoring were excluded, just like papers focusing on non-IT-related methods from the fields of electrical engineering, physics, and other fields not related to IT. Editorials, opinions, keynotes, abstracts, tutorial summaries, position papers, panel discussions, technical reports and posters were also excluded. The duplicated publications because of different versions were eliminated from consideration, with only the most recent version being included. Since this SLR considers peer-reviewed academic papers, grey literature was not covered in this research.
Inclusion and exclusion criteria are summarised below. For a paper to be included, all inclusion criteria described in Section III-B1 must be met. On the other hand, a paper is excluded if it matches any of the exclusion criteria from Section III-B1.
Inclusion Criteria
@@
  • Studies in the domain of cyber security in the maritime sector with a focus on security monitoring and intrusion detection.
  • Studies that propose novel cyber security monitoring and intrusion detection algorithms and technologies, and provide details about their application in the maritime domain.
Exclusion Criteria
  • Studies that do not explicitly focus on cyber security monitoring and intrusion detection.
@@
It should be pointed out that due to aforementioned criteria, a number of seemingly relevant maritime cyber security papers were excluded, because they did not focus on cyber security monitoring by contributing novel algorithms and technologies in this area. To provide a few examples of excluded papers, [27, 28] described maritime testbeds in the context of cyber attack tools and vulnerability testing without proposing new security monitoring approaches for detecting cyber attacks. In [29], a survey was conducted among maritime SOC experts to identify challenges in maritime SOCs without proposing novel security monitoring methods for addressing these challenges. In other words, the inclusion and exclusion criteria set the focus of the current study specifically to maritime cyber security monitoring papers with original algorithmical and technological contributions.
@@ III-B2 Search Strategy
Search for Related Publications
@@
Keyword Selection
@@
The selection of keywords heavily influences the results of the search. During the search phase, several keywords and their combinations were tried on search engines for electronic libraries. All the search engines that we used supported the creation of search queries with Boolean operators (AND, OR and NOT), which helped to narrow the focus of the search and limit the number of hits. Also, the engines allowed to apply the search expression on the title, abstract, keywords, metadata, etc. of the papers. The time frame of the search was also definable, and the wildcard characters (such as asterisk (*) and question mark (?)) allowed further flexibility during the search.
@@
In the first phase of keyword selection, the keywords related to the maritime sector, like “marine”, “ship”, and “navy” were selected to define the sector. These keywords combined with the OR-operator led to too many generic papers about marine systems. Then, the focus was narrowed by cyber security monitoring-specific keywords and phrases, like “monitoring”, “intrusion detection”, “IDS”, “attack detection”, etc. The queries which combined these keywords yielded many publications, but the majority of them remained out of scope. By adding “cyber” to these keywords, the focus was narrowed down further.
Automated Search on Bibliographical Sources
@@
During the initial stage of the automated search on bibliographical sources several trials were applied, using different combinations of keywords according to the main objectives of the review, with the aim to construct an optimal keyword set. The final keyword set was then used to construct the search query to be executed in various search engines of electronic libraries. The following electronic libraries were used to identify the publications relevant for the study: ACM Digital Library, IEEE Xplore, Elsevier Scopus, and Web of Science. Relevant publications were searched with the following query from all libraries: cyber AND (monitoring OR “intrusion detection” OR “anomaly detection” OR “attack detection” OR ids) AND (maritime OR navy OR ship OR marine OR sea OR ais)
Paper Selection Process
@@
The automated search from all libraries resulted in over 1000 papers. After the library search, the four co-authors processed one library each in the form of independent work, and the findings were added to a shared database that helped to indicate the duplicated items.
The next step in the selection of relevant papers was the analysis of the title and the abstract. If a paper was irrelevant based on the title and/or the abstract, it was excluded (In case of doubt, the researcher read the full text). If the title and abstract were in the scope, the paper became a candidate.
@@
If a paper remained relevant, its main properties were noted to support our findings. After the full-text reading of the candidates, we had a set of papers conforming to the inclusion criteria, and the entire procedure yielded 32 publications for a detailed analysis. At this point, the bibliometric analysis was written, and the methodology for analysis was developed.